OpenStudy (anonymous):

One number is two more than another number. If 4 times the larger number is subtracted from 5 times the smaller,the result is 10. Find both numbers.

OpenStudy (anonymous):

It says One number is it NEVER says another number is so another number gets the x another number = x one number = x + 2 which is the larger number "another" or "one"?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Which is the larger number another number which is x or one number which is x + 2

OpenStudy (anonymous):

i got Let X =smaller number and X=X=11 EQUALS the larger number

OpenStudy (anonymous):

5x - 4(x+2) = 10 5x - 4x - 8 = 10 x - 8 = 10 x = 18 18 and 20